# Adaptive Learning Guided by Bias-Noise-Alignment Diagnostics

### v1

@@ 3.5 Design Principles and Interpretation
The proposed bias-noise-alignment decomposition provides a compact and interpretable summary of learning dynamics where the bias component captures persistent error drift and signals the need to attenuate update magnitude when systematic overshoot is detected, the noise component captures stochastic variability and motivates conservative adaptation when feedback is unreliable, and the alignment component captures directional repetition and enables correction of oscillatory or resonant update behavior. Crucially, these diagnostics are orthogonal to model architecture, objective function, and learning paradigm. They rely only on scalar error signals and first-order update information, making them computationally lightweight and broadly applicable. Rather than replacing gradients or existing stabilization mechanisms, the diagnostics act as a higher-level control layer that modulates learning behavior based on the structure of the error signal. In the following sections, we show how this diagnostic backbone can be instantiated in supervised optimization, reinforcement learning, and meta-learning, yielding adaptive algorithms with improved stability, interpretability, and robustness under nonstationary conditions.
4 Diagnostic-Driven Supervised Optimization: HSAO
@@
for some constant $c_{1}>0$, indicating that diagnostic modulation preserves descent while improving robustness to nonstationary error dynamics. Algorithm 1 summarizes a single iteration of the proposed diagnostic-driven supervised optimizer.
@@
5 Diagnostic-Driven Reinforcement Learning: HED-RL
We next extend the proposed diagnostic framework to reinforcement learning and term the resulting method Hybrid Error-Diagnostic Reinforcement Learning (HED-RL). In RL, learning dynamics are governed by TD errors rather than supervised losses, introducing additional challenges due to bootstrapping, delayed rewards, and nonstationary data distributions induced by policy updates. We show that bias–noise–alignment diagnostics naturally generalize to TD-error signals and can be used to stabilize actor–critic learning through adaptive gating of critic updates, policy updates, and exploration. Consider a standard actor–critic setting with policy $\pi_{\theta}(a\mid s)$ and value function $V_{\psi}(s)$. The one-step TD error at time $t$ is
@@ 5 Diagnostic-Driven Reinforcement Learning: HED-RL
where $r_{t}$ is the reward and $\gamma\in(0,1)$ is the discount factor. The TD error serves as the primary learning signal for both the critic and the policy (via advantage estimates). Analogous to the supervised case, we focus on the temporal structure of TD errors and track exponential moving averages (EMAs) capturing bias and variability:
@@
which distinguish between regimes dominated by systematic TD bias such as the model mismatch or policy-induced drift and regimes dominated by stochastic variability such as the exploration noise or partial observability. We incorporate the TD-error diagnostics into actor–critic learning by modulating effective step sizes for the critic and the policy. The value function parameters $\psi$ are updated using a gated TD step:
@@
Since the policy gate satisfies $0<(1+k_{b}\rho^{\text{bias}}_{t})^{-1}\leq 1$, the effective policy step size is upper bounded by $\alpha_{\pi}$. Assuming bounded advantages and gradients, the policy update magnitude remains bounded at each iteration. The critic gate similarly attenuates updates when TD-error variance increases, reducing the impact of noisy bootstrap targets. Under standard assumptions of bounded rewards and contraction of the Bellman operator in expectation, the gated critic update preserves stability and improves robustness to variance spikes. Together, bias-based policy gating, noise-based critic gating, and diagnostic-driven entropy scheduling provide a coherent mechanism for stabilizing actor–critic learning; rather than relying on fixed heuristics, HED-RL adapts learning behavior based on interpretable TD-error structure. Algorithm 2 summarizes the HED-RL procedure for actor–critic learning.
Algorithm 2 HED-RL: Diagnostic-Driven Actor–Critic Learning
@@
6 Diagnostic-Conditioned Meta-Learning: MLLP
We finally extend the diagnostic-driven learning framework to meta-learning in the form of a Meta-Learned Learning Policy (MLLP), where the objective is to enable rapid and stable adaptation across a distribution of tasks. While meta-learning and learned optimizers have demonstrated impressive data efficiency, most existing approaches condition updates primarily on raw gradient information or opaque internal states. In contrast, we show that incorporating bias–noise–alignment diagnostics as explicit conditioning signals yields learned update rules that are both interpretable and robust under task heterogeneity.
@@ 7 Empirical Validation
This section provides illustrative empirical evidence to validate the behavior and interpretability of the proposed diagnostic-driven learning framework. The goal of these experiments is not to benchmark task-level performance or establish state-of-the-art results, but rather to demonstrate that (i) the proposed bias–noise–alignment diagnostics evolve in accordance with their theoretical interpretation, (ii) diagnostic-driven adaptation modulates learning dynamics in a stable and principled manner under nonstationarity, and (iii) the same diagnostic backbone generalizes consistently across supervised learning, reinforcement learning, and meta-learning paradigms.
Across all experiments, we focus on learning-dynamics indicators such as loss evolution, temporal-difference error statistics, update magnitudes, and diagnostic traces, rather than final task accuracy or control performance. All plots report representative runs and are intended to illustrate qualitative trends.
@@ 7.1 Supervised Learning Diagnostics
We discuss the behavior of the diagnostic-driven supervised optimizer (HSAO) in nonstationary supervised learning settings. The objective is to examine how bias, noise, and alignment diagnostics evolve during training and how they influence learning-rate modulation and update behavior. In such settings, periods of systematic drift in the error signal are reflected in sustained bias increases, which trigger attenuation of the effective learning rate. Stochastic perturbations manifest as spikes in the noise diagnostic, leading to more conservative updates. Alignment diagnostics rise during repeated directional excitation, correlating with overshoot-prone update behavior. These trends are consistent with the intended interpretation of the diagnostics and demonstrate how HSAO adapts learning dynamics based on error structure rather than instantaneous gradient information. Comprehensive empirical validation for supervised battery temperature estimation tasks is deferred to a companion experimental study.
@@ 7.2 Reinforcement Learning Diagnostics
We next illustrate the diagnostic behavior of the proposed Hybrid Error-Diagnostic Reinforcement Learning (HED-RL) framework in actor–critic training. Rather than focusing on episodic returns or task success rates, we analyze the temporal evolution of temporal-difference (TD) error diagnostics and their influence on critic updates, policy updates, and entropy regulation.
@@
Figure 1: Entropy coefficient during training. Baseline PPO uses a fixed coefficient, while HED-RL adapts it based on TD-error bias/noise diagnostics.
@@ 7.3 Meta-Learning Diagnostics
Finally, we discuss the role of diagnostic conditioning in the meta-learned learning policy (MLLP). In few-shot adaptation settings with heterogeneous task distributions, diagnostic signals provide explicit indicators of error reliability and curvature, enabling the learned optimizer to regulate update magnitudes across tasks. By conditioning inner-loop updates on bias and noise diagnostics rather than solely on gradient information, MLLP promotes smoother and more stable adaptation behavior under task-level uncertainty. Comprehensive empirical validation of diagnostic-conditioned meta-learning in battery-centric applications is deferred to a companion experimental study.
@@ 7.4 Diagnostic Ablation Analysis
These illustrative results support the central premise of this work: bias–noise–alignment diagnostics provide interpretable, early indicators of learning instability and enable principled modulation of learning dynamics across diverse learning paradigms.
@@ 8 Unified Perspective, Implications, and Limitations
This work proposes a diagnostic-driven perspective on adaptive learning that treats error evolution as a first-class object rather than an incidental byproduct of gradient-based optimization. The experiments are intentionally designed to validate diagnostic behavior and stability mechanisms, rather than to maximize task-specific performance. By decomposing error dynamics into bias, noise, and alignment components, we introduce a unifying framework that spans supervised learning, reinforcement learning, and meta-learning, while remaining low-overhead, interpretable, and task-agnostic.
In particular, the diagnostics are derived from finite-window EMAs and can lag abrupt regime changes. Across learning paradigms, adaptation decisions are traditionally driven by instantaneous gradient information or heuristic schedules. In contrast, the proposed framework leverages the temporal structure of error signals to regulate learning behavior. Bias captures persistent drift indicating systematic mismatch between model updates and the evolving objective. Noise captures stochastic variability that undermines the reliability of feedback. Alignment captures directional repetition that can lead to oscillation or resonance. These diagnostics provide complementary information that is largely orthogonal to gradient magnitude or curvature, and they are computed from simple statistics of scalar error signals without requiring higher-order derivatives, model internals, or task-specific structure. As a result, the same diagnostic backbone naturally applies to supervised losses, temporal-difference errors in reinforcement learning, and task-level losses in meta-learning.
@@
Despite these limitations, the proposed framework suggests a broader design principle for future learning systems: reliable adaptation should be driven not only by gradients, but also by the structure and reliability of error signals. This perspective opens avenues for integrating diagnostic-driven adaptation with probabilistic learning, uncertainty estimation, and system-level monitoring, and provides a foundation for application-specific studies in safety-critical and nonstationary domains.
@@ 9 Conclusion
This paper introduced a diagnostic-driven framework for adaptive learning based on a principled decomposition of error dynamics into bias, noise, and alignment components. By explicitly modeling the temporal evolution of error signals, the proposed framework provides a lightweight and interpretable control layer that augments gradient-based learning without altering model architectures or objective functions. We demonstrated that the same diagnostic backbone naturally extends across supervised optimization, actor–critic reinforcement learning, and meta-learning, yielding diagnostic-driven instantiations with bounded updates and improved stability under standard smoothness assumptions. Unlike conventional adaptive methods that rely primarily on instantaneous gradient statistics, the proposed approach regulates learning behavior based on the reliability and structure of the error signal itself. More broadly, this work advocates a shift in perspective: reliable adaptation should be guided not only by gradients, but also by interpretable diagnostics of error evolution. By elevating error dynamics to a first-class object, the framework bridges ideas from optimization, control, and meta-learning, and establishes a foundation for the principled design of adaptive learning systems in nonstationary environments. We hope this perspective will stimulate further theoretical and empirical research on diagnostic-driven learning and its integration with uncertainty-aware and control-oriented learning frameworks.
@@ Acknowledgments
The author gratefully acknowledges Trihaan Samanta, son of Akash Samanta, whose everyday observations of human learning behavior helped inspire the intuition underlying this work: that systematic bias and stochastic noise can be disentangled and adaptively regulated during learning, much as humans naturally refine their actions through experience.
@@ Appendix A.1 Assumptions
We adopt the following standard assumptions commonly used in stochastic optimization and reinforcement learning:
@@
  • Rewards and temporal-difference errors are bounded.
@@ Appendix A.5 Bounded Learned Updates in MLLP
Assume the learned optimizer outputs bounded step sizes and gating coefficients. Then the inner-loop updates in (31) are bounded.

### v2

@@ 3.5 Design Principles and Interpretation
The proposed bias-noise-alignment decomposition provides a compact and interpretable summary of learning dynamics where the bias component captures persistent error drift and signals the need to attenuate update magnitude when systematic overshoot is detected, the noise component captures stochastic variability and motivates conservative adaptation when feedback is unreliable, and the alignment component captures directional repetition and enables correction of oscillatory or resonant update behavior. Crucially, these diagnostics are orthogonal to model architecture, objective function, and learning paradigm. They rely only on scalar error signals and first-order update information, making them computationally lightweight and broadly applicable. Rather than replacing gradients or existing stabilization mechanisms, the diagnostics act as a higher-level control layer that modulates learning behavior based on the structure of the error signal. In the following sections, we show how this diagnostic backbone can be instantiated in supervised optimization, RL, and meta-learning, yielding adaptive algorithms with improved stability, interpretability, and robustness under nonstationary conditions.
4 Diagnostic-Driven Supervised Optimization
@@
for some constant $c_{1}>0$, indicating that diagnostic modulation preserves descent while improving robustness to nonstationary error dynamics. Algorithm 1 summarizes a single iteration of the proposed diagnostic-driven supervised optimizer. Before applying Adam-style moment updates (10)–(11), HSAO updates the bias/noise EMAs (2)–(5) from the incremental error signal and then computes the gated step size.
@@
5 Diagnostic-Driven Reinforcement Learning
We next extend the proposed diagnostic framework to RL and term the resulting method Hybrid Error-Diagnostic Reinforcement Learning (HED-RL). In RL, learning dynamics are governed by TD errors rather than supervised losses, introducing additional challenges due to bootstrapping, delayed rewards, and nonstationary data distributions induced by policy updates. We show that bias-noise-alignment diagnostics naturally generalize to TD-error signals and can be used to stabilize actor-critic learning through adaptive gating of critic updates, policy updates, and exploration. Consider a standard actor-critic setting with policy $\pi_{\theta}(a\mid s)$ and value function $V_{\psi}(s)$. The one-step TD error at time $t$ is
@@ 5 Diagnostic-Driven Reinforcement Learning
where $r_{t}$ is the reward and $\gamma\in(0,1)$ is the discount factor. The TD error serves as the primary learning signal for both the critic and the policy via advantage estimates. Analogous to the supervised case, we focus on the temporal structure of TD errors and track EMAs capturing bias and variability:
@@
which distinguish between regimes dominated by systematic TD bias such as the model mismatch or policy-induced drift and regimes dominated by stochastic variability such as the exploration noise or partial observability. We incorporate the TD-error diagnostics into actor-critic learning by modulating effective step sizes for the critic and the policy. The value function parameters $\psi$ are updated using a gated TD step:
@@
Since the policy gate satisfies $0<(1+k_{b}\rho^{\text{bias}}_{t})^{-1}\leq 1$, the effective policy step size is upper bounded by $\alpha_{\pi}$. Assuming bounded advantages and gradients, the policy update magnitude remains bounded at each iteration. The critic gate similarly attenuates updates when TD-error variance increases, reducing the impact of noisy bootstrap targets. Under standard assumptions of bounded rewards and contraction of the Bellman operator in expectation, the gated critic update preserves stability and improves robustness to variance spikes. Together, bias-based policy gating, noise-based critic gating, and diagnostic-driven entropy scheduling provide a coherent mechanism for stabilizing actor-critic learning; rather than relying on fixed heuristics, HED-RL adapts learning behavior based on interpretable TD-error structure. Algorithm 2 summarizes the HED-RL procedure for actor-critic learning.
Algorithm 2 HED-RL: Diagnostic-Driven actor-critic Learning
@@
6 Diagnostic-Conditioned Meta-Learning
We finally extend the diagnostic-driven learning framework to meta-learning in the form of a Meta-Learned Learning Policy (MLLP), where the objective is to enable rapid and stable adaptation across a distribution of tasks. While meta-learning and learned optimizers have demonstrated impressive data efficiency, most existing approaches condition updates primarily on raw gradient information or opaque internal states. In contrast, we show that incorporating bias-noise-alignment diagnostics as explicit conditioning signals yields learned update rules that are both interpretable and robust under task heterogeneity.
@@ 7 Empirical Validation
This section provides illustrative empirical evidence to validate the behavior and interpretability of the proposed diagnostic-driven learning framework. The goal of these experiments is not to benchmark task-level performance or establish state-of-the-art results, but rather to demonstrate that (i) the proposed bias-noise-alignment diagnostics evolve in accordance with their theoretical interpretation, (ii) diagnostic-driven adaptation modulates learning dynamics in a stable and principled manner under nonstationarity, and (iii) the same diagnostic backbone generalizes consistently across supervised learning, RL, and meta-learning paradigms.
Across all experiments, we focus on learning-dynamics indicators such as loss evolution, TD error statistics, update magnitudes, and diagnostic traces, rather than final task accuracy or control performance. All plots report representative runs and are intended to illustrate qualitative trends.
@@ 7.1 Supervised Learning Diagnostics
We discuss the behavior of HSAO in nonstationary supervised learning settings. The objective is to examine how bias, noise, and alignment diagnostics evolve during training and how they influence learning-rate modulation and update behavior. In such settings, periods of systematic drift in the error signal are reflected in sustained bias increases, which trigger attenuation of the effective learning rate. Stochastic perturbations manifest as spikes in the noise diagnostic, leading to more conservative updates. Alignment diagnostics rise during repeated directional excitation, correlating with overshoot-prone update behavior. These trends are consistent with the intended interpretation of the diagnostics and demonstrate how HSAO adapts learning dynamics based on error structure rather than instantaneous gradient information. Comprehensive empirical validation for supervised battery temperature estimation tasks is deferred to a companion experimental study.
@@ 7.2 Reinforcement Learning Diagnostics
We next illustrate the diagnostic behavior of the proposed HED-RL framework in actor-critic training. Rather than focusing on episodic returns or task success rates, we analyze the temporal evolution of TD error diagnostics and their influence on critic updates, policy updates, and entropy regulation.
@@
Figure 1: Entropy coefficient during training. Baseline PPO uses a fixed coefficient, while HED-RL adapts it based on TD-error bias/noise diagnostics (representative run).
@@ 7.3 Meta-Learning Diagnostics
Finally, we discuss the role of diagnostic conditioning in the MLLP. In few-shot adaptation settings with heterogeneous task distributions, diagnostic signals provide explicit indicators of error reliability and curvature, enabling the learned optimizer to regulate update magnitudes across tasks. By conditioning inner-loop updates on bias and noise diagnostics rather than solely on gradient information, MLLP promotes smoother and more stable adaptation behavior under task-level uncertainty. Comprehensive empirical validation of diagnostic-conditioned meta-learning in battery-centric applications is deferred to a companion experimental study.
@@ 7.4 Diagnostic Ablation Analysis
These illustrative results support the central premise of this work: bias-noise-alignment diagnostics provide interpretable, early indicators of learning instability and enable principled modulation of learning dynamics across diverse learning paradigms.
@@ 8 Unified Perspective, Implications, and Limitations
This work proposes a diagnostic-driven perspective on adaptive learning that treats error evolution as a first-class object rather than an incidental byproduct of gradient-based optimization. The experiments are intentionally designed to validate diagnostic behavior and stability mechanisms, rather than to maximize task-specific performance. By decomposing error dynamics into bias, noise, and alignment components, we introduce a unifying framework that spans supervised learning, RL, and meta-learning, while remaining low-overhead, interpretable, and task-agnostic.
In particular, the diagnostics are derived from finite-window EMAs and can lag abrupt regime changes. Across learning paradigms, adaptation decisions are traditionally driven by instantaneous gradient information or heuristic schedules. In contrast, the proposed framework leverages the temporal structure of error signals to regulate learning behavior. Bias captures persistent drift indicating systematic mismatch between model updates and the evolving objective [hu2021bias]. Noise captures stochastic variability that undermines the reliability of feedback [wojtowytsch2023stochastic]. Alignment captures directional repetition that can lead to oscillation or resonance. These diagnostics provide complementary information that is largely orthogonal to gradient magnitude or curvature, and they are computed from simple statistics of scalar error signals without requiring higher-order derivatives, model internals, or task-specific structure. As a result, the same diagnostic backbone naturally applies to supervised losses, TD errors in RL, and task-level losses in meta-learning.
@@
Despite these limitations, the proposed framework suggests a broader design principle for future learning systems: reliable adaptation should be driven not only by gradients, but also by the structure and reliability of error signals. This perspective opens avenues for integrating diagnostic-driven adaptation with probabilistic learning, uncertainty estimation, and system-level monitoring, and provides a foundation for application-specific studies in safety-critical and nonstationary domains [karnehm2025core, 11317983].
@@ 9 Conclusion
This paper introduced a diagnostic-driven framework for adaptive learning based on a principled decomposition of error dynamics into bias, noise, and alignment components. By explicitly modeling the temporal evolution of error signals, the proposed framework provides a lightweight and interpretable control layer that augments gradient-based learning without altering model architectures or objective functions. We demonstrated that the same diagnostic backbone naturally extends across supervised optimization, actor-critic RL, and meta-learning, yielding diagnostic-driven instantiations with bounded updates and improved stability under standard smoothness assumptions. Unlike conventional adaptive methods that rely primarily on instantaneous gradient statistics, the proposed approach regulates learning behavior based on the reliability and structure of the error signal itself. More broadly, this work advocates a shift in perspective: reliable adaptation should be guided not only by gradients, but also by interpretable diagnostics of error evolution. By elevating error dynamics to a first-class object, the framework bridges ideas from optimization, control, and meta-learning, and establishes a foundation for the principled design of adaptive learning systems in nonstationary environments. We hope this perspective will stimulate further theoretical and empirical research on diagnostic-driven learning and its integration with uncertainty-aware and control-oriented learning frameworks.
@@ Acknowledgments
The authors gratefully acknowledge Trihaan Samanta, whose everyday observations of human learning behavior helped inspire the intuition underlying this work, that systematic bias and stochastic noise can be disentangled and adaptively regulated during learning through experience.
Funding
This research received no external funding.
@@ Appendix A.1 Assumptions
We adopt the following standard assumptions commonly used in stochastic optimization and RL:
@@
  • Rewards and TD errors are bounded.
@@ Appendix A.5 Bounded Learned Updates in MLLP
Assume the learned optimizer outputs bounded step sizes and gating coefficients. Then the inner-loop updates in (31)–(34) are bounded.